# itemsep=1pt,topsep=2pt,parsep=1pt,partopsep=1pt

### v2

@@
Abstract.
In this paper, we study a disordered pinning model induced by a random walk whose increments have a finite fourth moment and vanishing first and third moments. It is known that this model is marginally relevant, and moreover, it undergoes a phase transition in an intermediate disorder regime. We show that, in the critical window, the point-to-point partition functions converge to a unique limiting random measure, which we call the critical disordered pinning measure. We also obtain an analogous result for a continuous counterpart to the pinning model, which is closely related to two other models: one is a critical stochastic Volterra equation that gives rise to a rough volatility model, and the other is a critical stochastic heat equation with multiplicative noise that is white in time and delta in space.
MSC 2010. Primary: 82B44, Secondary: 35R60, 60H15, 91G20.
Keywords. Marginal relevance, critical temperature, disordered pinning model, directed polymer, stochastic heat equation, rough volatility model.
1. Introduction
1.1. Background
In this paper, we study the disordered pinning model (abbreviated as the pinning model hereafter), which is also known as a model of pinning on a defect line (see, e.g., the monographs of Giacomin [Gia07, Gia11] and den Hollander [dH09] for a comprehensive introduction). This model holds significance in statistical mechanics, serving as a disordered system that is both interesting enough to have a phase transition and simple enough for extensive mathematical analysis.
@@ 1.1. Background
(1.1) ${\mathbbm{E}}[\omega_{0}]=0,\quad{\rm Var}(\omega_{0})=1,\quad\text{and}\quad\lambda(\beta):=\log{\mathbbm{E}}[\exp(\beta\omega_{0})]<+\infty,\quad\forall\beta\in(-\beta_{0},\beta_{0}),$
@@
(1.2) $\frac{{\mathrm{d}}{\mathbf{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbf{P}}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big{(}\sum_{n=1}^{N}(\beta\omega_{n}+h-\lambda(\beta))\mathbbm{1}_{\{n\in\tau\}}\Big{)},$
@@
(1.3) $Z_{N,\beta}^{\omega,h}:={\mathbf{E}}\Big{[}\exp\Big{(}\sum_{n=1}^{N}(\beta\omega_{n}+h-\lambda(\beta))\mathbbm{1}_{\{n\in\tau\}}\Big{)}\Big{]}$
@@
A particularly interesting setup is to assume the renewal $\tau$ is recurrent, i.e., ${\mathbf{P}}(\tau_{1}<\infty)=1$, and
(1.4) $K(n):={\mathbf{P}}(\tau_{1}=n)=\frac{L(n)}{n^{1+\alpha}}$
for some $\alpha\geq 0$ and slowly varying function $L(\cdot)$. Examples of slowly varying functions include constant functions and logarithmic functions (see [BGT89] for details). For different values of $\alpha$, determining whether the disordered measure ${\mathbf{P}}_{N,\beta}^{\omega,h}$ significantly differs from the pure measure ${\mathbf{P}}_{N}^{h}$ for arbitrarily small $\beta>0$ becomes a subtle problem, which we will elucidate below.
@@
The existence and non-randomness of the limit are proven in [Gia11]. Moreover, it shows that for any $\beta\geq 0$ fixed, $F(\beta,h)$ is non-negative, convex, and non-decreasing as a function of $h$. Thus, the system undergoes a phase transition in $h$ at
(1.6) $h_{c}(\beta):=\inf\{h:F(\beta,h)>0\}.$
The phase $\{(\beta,h):F(\beta,h)>0\}$ is called the localized regime, wherein, as $N\to\infty$, the number of renewal times under ${\mathbf{P}}_{N,\beta}^{\omega,h}$ is proportional to $N$. This signifies that the energy gain dominates the entropy cost for the renewal process to visit the origin with atypical high frequency. Conversely, the phase $\{(\beta,h):F(\beta,h)=0\}$ is called the delocalized regime, where the number of renewal times is $o(N)$, as the energy gain at renewal returns is less attractive to the renewal process.
@@
(1.7) ${\mathbbm{E}}\big{[}\log Z_{N,\beta}^{\omega,h}\big{]}\leq\log{\mathbbm{E}}\big{[}Z_{N,\beta}^{\omega,h}\big{]}\quad\text{with}\quad{\mathbbm{E}}\big{[}Z_{N,\beta}^{\omega,h}\big{]}={\mathbf{E}}\Big{[}\exp\Big{(}{\sum_{n=0}^{N}h\mathbbm{1}_{\{n\in\tau\}}}\Big{)}\Big{]}=Z_{N}^{h},$
where $Z_{N}^{h}$ is the partition function of the pure pinning model ($\beta=0$). Note that the pure free energy $F(h)=F(0,h)=\lim\limits_{N\to\infty}\frac{1}{N}\log Z_{N}^{h}$ is deterministic and has an explicit form. It is well established (see [Gia11]) that $F(h)=0$ if and only if $h\leq 0$, i.e., $h_{c}:=\inf\{h\in{\mathbbm{R}}:F(h)>0\}=0$, and moreover,
(1.8) $\lim\limits_{h\downarrow h_{c}}\frac{\log F(h)}{\log(h-h_{c})}=1\vee\alpha^{-1}.$
@@
The key question for the pinning model is: does arbitrarily small disorder ($\beta>0$) qualitatively change the critical behavior of the system? In particular, one may ask whether for arbitrarily small $\beta>0$, the critical exponent for $F(\beta,h)$ differs from $1\vee\alpha^{-1}$ that appears in (1.8) for the pure model. It turns out that this can be equivalently characterised in terms of critical point shift, namely whether $h_{c}(\beta)>h_{c}(0)=0$ for arbitrarily small $\beta>0$, or there exists some $\beta^{\ast}>0$ such that $h_{c}(\beta)=0$ for $\beta\in[0,\beta^{\ast}]$. The answer hinges on the exponent $\alpha$ in (1.4). It is predicted by the Harris criterion [Har74] that the disorder is relevant (resp. irrelevant) if $\alpha>1/2$ (resp. $\alpha<1/2$). Indeed, it is now known that
@@
$\bullet$ For $\alpha<\frac{1}{2}$, by [Ale08, Ton08, Lac10] the critical exponents for both the disordered model and pure model are the same, and $h_{c}(\beta)\equiv 0=h_{c}$ for $\beta>0$ small enough. This regime is known as the disorder irrelevant regime.
@@
$\bullet$ For $\alpha>\frac{1}{2}$, it was shown by Giacomin and Toninelli [GT06] that the critical exponent for the disordered model is at least 2, which is strictly larger than $1\vee\alpha^{-1}$ in (1.8). Furthermore, [AZ09, DGLT09] established that $h_{c}(\beta)>0$ for any $\beta>0$, and later [${}^{+}$14, CTT17] obtained the sharp asymptotics for $h_{c}(\beta)$ as $\beta\to 0$. This is called the disorder relevant regime.
@@
$\bullet$ For $\alpha=\frac{1}{2}$, the situation becomes intricate, and it is out of the scope of the Harris criterion. A crucial quantity is $R_{N}:=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}$, the expected number of overlaps for two i.i.d. copies of the renewal process $\tau$. It is shown in [BL18] that $h_{c}(\beta)>0$ for any $\beta>0$ (and hence the disorder is relevant) if and only if $R_{N}\sim\sum_{n=1}^{N}1/(nL(n)^{2})\to+\infty$. In this scenario, it is called the marginally relevant regime, and [BL18] provides the sharp asymptotics for $h_{c}(\beta)$ as $\beta\to 0$.
@@
In a nutshell, for $\alpha\geq\frac{1}{2}$ and when the disorder is relevant, we observe significantly different behavior of the pinning model near the critical curve $h=h_{c}(\beta)$ at $(\beta,h_{c}(\beta))=(0,0)$. The disordered system and the pure system belong to two universality classes (cf. [CSZ16, Section 1.3]): it is believed that there are different universal scaling limits for the two systems, where neither the disorder distribution nor the details of the renewal process are relevant (as long as some mild conditions hold, e.g., (1.1) holds and the exponent $\alpha$ is fixed).
@@
It is then interesting to study the interpolation of the two universalities. In view of the partition function, as per (1.5), in the localized regime, $Z_{N,\beta}^{\omega,h}$ exhibits exponential growth, determined by the exponent $NF(\beta,h)$. This exponent increases as $N$ tends to infinity and decreases as both $\beta$ and $h$ approach 0. Conversely, in the delocalized regime, $Z_{N,\beta}^{\omega,h}$ demonstrates sub-exponential growth, with $Z_{N,0}^{\omega,0}\equiv 1$. A natural question about interpolating is: if we systematically tune down $\beta:=\beta_{N}\downarrow 0$ and $h:=h_{N}\downarrow 0$ at specific rates as $N\to\infty$, can we ascertain a non-trivial limit for $Z_{N,\beta_{N}}^{\omega,h_{N}}$? The affirmative answer implies the existence of a limit, termed the weak coupling limit, situating the model within the intermediate disorder regime. The inception of exploring the intermediate disorder regime traces back to [AKQ14], focusing on the $1+1$ directed polymer model. Subsequent investigations into weak coupling limits include disorder relevant models such as the 2d random field Ising model, the pinning model with $\alpha>1/2$ in [CSZ16], and marginally relevant models such as the 2d stochastic heat equation, the $1+2$ directed polymer model, and the pinning model with $\alpha=1/2$ in [CSZ17].
@@
For the pinning model with $\alpha>\frac{1}{2}$, a non-trivial weak coupling limit emerges through the choice of $\beta_{N}=\widehat{\beta}{\mathcal{P}}_{1}^{\alpha}(N)$ and $h_{N}=\widehat{h}{\mathcal{P}}_{2}^{\alpha}(N)$ where $\widehat{\beta}>0,\widehat{h}\in{\mathbbm{R}}$, and ${\mathcal{P}}_{1}^{\alpha}(\cdot)$ and ${\mathcal{P}}_{2}^{\alpha}(\cdot)$, both vanishing at infinity, are two regularly varying functions whose exponents depending on $\alpha$ (see [CSZ16] for details). However, the scenario becomes more sophisticated in the marginal case $\alpha=\frac{1}{2}$, as illustrated in [CSZ17]. The appropriate choice is $h=0$ and $\beta_{N}=\widehat{\beta}/\sqrt{R_{N}}$, where $R_{N}=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}\to+\infty$ as above and is slowly varying at infinity. Furthermore, the constant $\widehat{\beta}$ plays a pivotal role. If $\widehat{\beta}\in(0,1)$, then the weak coupling limit for $Z_{N,\beta_{N}}^{\omega,0}$ is log-normal, characterizing a subcritical phase. For the critical phase $\widehat{\beta}=1$ and the supercritical phase $\widehat{\beta}>1$, $Z_{N,\beta_{N}}^{\omega,0}\to 0$ in probability as $N\to\infty$.
@@
While the behavior of the partition function in the supercritical phase remains mysterious, a recent breakthrough by Caravenna, Sun, and Zygouras in [CSZ23] concerning the marginally relevant directed polymer in two dimension sheds light on the critical phase. They viewed the random field of point-to-point partition functions as a process of random measures on ${\mathbbm{R}}^{2}\times{\mathbbm{R}}^{2}$, and established the existence of a universal scaling limit known as the critical 2d stochastic heat flow. Moreover, the limit can be interpreted as the solution of the stochastic heat equation in the critical dimension 2 and in the critical window, which lies beyond existing solution theories via regularity structures by Hairer [Hai14] and paracontrolled distributions by Gubinelli et. al. [GIP15] for subcritical singular SPDEs. Motivated by their work, our study in this paper is centered on the critical phase of the marginally relevant pinning model (Subsection 1.2), as well as a rough volatility model characterised by the critical stochastic Volterra equation (Subsection 1.3).
1.2. The pinning model and weak coupling limit
@@
Instead of the most general setting (1.2), we focus on a specific class of pinning models associated to a random walk $S:=(S_{n})_{n\geq 0}$ on $\mathbb{Z}$ due to technical limitation. To be specific, our proof requires that the renewal process is associated to a Markovian process with sharp enough local limit theorems (see Remark 1.4 and Section 4). Noting that the times $S$ visits 00 give rise to a renewal process, it naturally induces a pinning model. Denote the probability law of $S$ by ${\mathbf{P}}^{x}$ if ${\mathbf{P}}^{x}(S_{0}=x)=1$, indicating that the starting point of $S$ is $x$. The expectation with respect to ${\mathbf{P}}^{x}$ is denoted by ${\mathbf{E}}^{x}$. When $x=0$, we omit the superscripts in ${\mathbf{P}}$ and ${\mathbf{E}}$ to lighten the notation. The distribution of the induced renewal process is then given by
(1.9) $K(n)={\mathbf{P}}(\tau_{1}=n)={\mathbf{P}}(S_{n}=0,S_{k}\neq 0,1\leq k\leq n-1)\quad\text{and}\quad{\mathbf{P}}(n\in\tau)={\mathbf{P}}(S_{n}=0).$
@@
(1.10) ${\mathbf{E}}[S_{1}]=0,\quad{\rm Var}(S_{1})=1,\quad{\mathbf{E}}[S_{1}^{3}]=0,\quad\text{and}\quad{\mathbf{E}}[|S_{1}|^{4}]<+\infty,$
where the assumptions on the higher moments are due to technical reasons (see Remark 1.4 below).
It is known (see [Uch11]) that for a random walk on ${\mathbbm{Z}}$ with mean 00 and finite variance, the induced renewal process $\tau$ satisfies $K(n)\sim 1/(\sqrt{2\pi}\hskip 2.0ptn^{3/2})$, which falls within the framework of the marginally relevant pinning model, since $\alpha=\frac{1}{2}$ in (1.4). Furthermore, ${\mathbf{P}}(S_{n}=0)\sim 1/\sqrt{2\pi n}$ by the local limit theorem, and
(1.11) $R_{N}=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}=\sum_{n=1}^{N}{\mathbf{P}}(S_{n}=0)^{2}\sim\frac{\log N}{2\pi}\to\infty,\quad\text{as~{}}N\to\infty.$
@@
(1.12) $\frac{{\mathrm{d}}{\mathbf{P}}_{N,\beta_{N}}^{\omega}}{{\mathrm{d}}{\mathbf{P}}}(S):=\frac{1}{Z_{N,\beta_{N}}^{\omega}}\exp\Big{(}\sum_{n=1}^{N}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big{)},$
@@
(1.13) $Z_{N,\beta_{N}}^{\omega}:={\mathbf{E}}\Big{[}\exp\Big{(}\sum_{n=1}^{N}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big{)}\Big{]}.$
Note that as in [CSZ17], here we set the external field $h\equiv 0$ to study the intermediate disorder regime, which is for simplicity and not essential. We are satisfied with this setting and will briefly discuss nonzero choices of $h$ later in Remark 2.4.
Our goal is to study the weak coupling limit of $Z_{N,\beta_{N}}^{\omega}$ at the critical temperature $\beta_{N,c}:=1/\sqrt{R_{N}}\sim\sqrt{2\pi}/\sqrt{\log N}$. Recall that $Z_{N,\beta_{N,c}}^{\omega}$ converges in probability to the trivial limit 0 as $N\to\infty$. The right way to study the weak coupling limit, in the same spirit as [CSZ23], is to introduce a random measure related to $Z_{N,\beta_{N,c}}^{\omega}$ and consider its scaling limit. To this end, for given starting and end times $M\leq K$, we introduce a point-to-point pinning partition function
(1.14) $Z_{M,K}^{\omega,\beta_{N}}:={\mathbf{E}}\Big{[}\exp\Big{(}\sum\limits_{n=M}^{K}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big{)}\mathbbm{1}_{\{S_{K}=0\}}\Big{|}S_{M}=0\Big{]}.$
@@
(1.15) ${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t):=\sqrt{N}Z_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}{\mathrm{d}}s{\mathrm{d}}t,$
@@
(1.16) $\sqrt{N}{\mathbbm{E}}\big{[}Z_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}\big{]}=\sqrt{N}{\mathbf{P}}\big{(}S_{\lfloor tN\rfloor-\lfloor sN\rfloor}\rfloor=0\big{)}\sim\frac{1}{\sqrt{2\pi(t-s)}}.$
We equip the space of locally finite measures on ${\mathbbm{R}}^{2}_{\leq}$ with the topology of vague convergence
(1.17) $\mu_{n}\to\mu\quad\Longleftrightarrow\quad\int\phi(s,t)\mu_{n}({\mathrm{d}}s,{\mathrm{d}}t)\overset{\rm d}{\longrightarrow}\int\phi(s,t)\mu({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\phi\in C_{c}({\mathbbm{R}}^{2}_{\leq}).$
@@
(1.18) $\zeta_{n}=\zeta_{n}^{(N)}:=e^{\beta_{N}\omega_{n}-\lambda(\beta_{N})}-1.$

### v3

# The critical disordered pinning measure

@@
Abstract
In this paper, we study a disordered pinning model induced by a random walk whose increments have a finite $(2+\kappa)$-th moment for some $\kappa>0$. It is known that this model is marginally relevant, and moreover, it undergoes a phase transition in an intermediate disorder regime. We show that, in the critical window, the point-to-point partition functions converge to a unique limiting random measure, which we call the critical disordered pinning measure. We also obtain an analogous result for a continuous counterpart to the pinning model, which is closely related to two other models: one is a critical stochastic Volterra equation that gives rise to a rough volatility model, and the other is a critical stochastic heat equation with multiplicative noise that is white in time and delta in space.
rough volatility model,
keywords:
[class=MSC]
keywords:
[…]
, and
1 Introduction
1.1 Background
In this paper, we study the disordered pinning model (abbreviated as the pinning model hereafter), which is also known as a model of pinning on a defect line (see, e.g., the monographs of Giacomin [21, 22] and den Hollander [19] for a comprehensive introduction). This model holds significance in statistical mechanics, serving as a disordered system that is both interesting enough to have a phase transition and simple enough for extensive mathematical analysis.
@@ 1.1 Background
${\mathbbm{E}}[\omega_{0}]=0,~~{\rm Var}(\omega_{0})=1,~~\text{and}~~\lambda(\beta):=\log{\mathbbm{E}}[\exp(\beta\omega_{0})]<+\infty,~~\forall\beta\in(-\beta_{0},\beta_{0}),$ (1.1)
@@
$\frac{{\mathrm{d}}{\mathbf{P}}_{N,\beta}^{\omega,h}}{{\mathrm{d}}{\mathbf{P}}}(\tau):=\frac{1}{Z_{N,\beta}^{\omega,h}}\exp\Big(\sum_{n=1}^{N}(\beta\omega_{n}+h-\lambda(\beta))\mathbbm{1}_{\{n\in\tau\}}\Big),$ (1.2)
@@
$Z_{N,\beta}^{\omega,h}:={\mathbf{E}}\Big[\exp\Big(\sum_{n=1}^{N}(\beta\omega_{n}+h-\lambda(\beta))\mathbbm{1}_{\{n\in\tau\}}\Big)\Big]$ (1.3)
@@
A particularly interesting setup is to assume the renewal $\tau$ is recurrent, i.e., ${\mathbf{P}}(\tau_{1}<\infty)=1$, and has a distribution
$K(n):={\mathbf{P}}(\tau_{1}=n)=\frac{L(n)}{n^{1+\alpha}}$ (1.4)
for some $\alpha\geq 0$ and some slowly varying function $L(\cdot)$. Examples of slowly varying functions include $L(x)\equiv C$ and $L(x)=\log x$ (see [10] for details). For different values of $\alpha$, determining whether the disordered measure ${\mathbf{P}}_{N,\beta}^{\omega,h}$ significantly differs from the pure measure ${\mathbf{P}}_{N}^{h}$ for arbitrarily small $\beta>0$ becomes a subtle problem, which we will elucidate below.
@@
The existence and non-randomness of the limit are proven in [22]. Moreover, it shows that for any $\beta\geq 0$ fixed, $F(\beta,h)$ is non-negative, convex, and non-decreasing as a function of $h$. Thus, the system undergoes a phase transition in $h$ at
$h_{c}(\beta):=\inf\{h:F(\beta,h)>0\}.$ (1.6)
The phase $\{(\beta,h):F(\beta,h)>0\}$ is called the localized regime, wherein, as $N\to\infty$, the number of renewal times under ${\mathbf{P}}_{N,\beta}^{\omega,h}$ is proportional to $N$. This signifies that the energy gain dominates the entropy cost for the renewal process to visit the origin with atypical high frequency. On the other hand, the phase $\{(\beta,h):F(\beta,h)=0\}$ is called the delocalized regime, where the number of renewal times is $o(N)$, as the energy gain at renewal returns is less attractive to the renewal process.
@@
${\mathbbm{E}}\big[\log Z_{N,\beta}^{\omega,h}\big]\leq\log{\mathbbm{E}}\big[Z_{N,\beta}^{\omega,h}\big]\quad\text{with}\quad{\mathbbm{E}}\big[Z_{N,\beta}^{\omega,h}\big]={\mathbf{E}}\Big[\exp\Big({\sum_{n=0}^{N}h\mathbbm{1}_{\{n\in\tau\}}}\Big)\Big]=Z_{N}^{h},$ (1.7)
where $Z_{N}^{h}$ is the partition function of the pure pinning model ($\beta=0$). Note that the pure free energy $F(h)=F(0,h)=\lim_{N\to\infty}\frac{1}{N}\log Z_{N}^{h}$ is deterministic and has an explicit form. It is well established (see [22]) that $F(h)=0$ if and only if $h\leq 0$, i.e., $h_{c}:=\inf\{h\in{\mathbbm{R}}:F(h)>0\}=0$, and moreover,
$\lim\limits_{h\downarrow h_{c}}\frac{\log F(h)}{\log(h-h_{c})}=1\vee\alpha^{-1}.$ (1.8)
@@
The key question for the pinning model is: does arbitrarily small disorder ($\beta>0$) qualitatively change the critical behavior of the system? In particular, one may ask whether for arbitrarily small $\beta>0$, the critical exponent for $F(\beta,h)$ differs from $1\vee\alpha^{-1}$ that appears in (1.8) for the pure model. It turns out that this can be equivalently characterised in terms of critical point shift, namely whether $h_{c}(\beta)>h_{c}(0)=0$ for arbitrarily small $\beta>0$, or there exists some $\beta^{\ast}>0$ such that $h_{c}(\beta)=0$ for $\beta\in[0,\beta^{\ast}]$. The answer hinges on the exponent $\alpha$ in (1.4). It is predicted by the Harris criterion [26] that the disorder is relevant (resp. irrelevant) if $\alpha>1/2$ (resp. $\alpha<1/2$). Indeed, it is now known that
@@
$\bullet$ For $\alpha<\frac{1}{2}$, by [2, 35, 29] the critical exponents for both the disordered model and the pure model are the same, and $h_{c}(\beta)\equiv 0=h_{c}$ for $\beta>0$ small enough. This regime is known as the disorder irrelevant regime.
@@
$\bullet$ For $\alpha>\frac{1}{2}$, it was shown by Giacomin and Toninelli [23] that the critical exponent for the disordered model is at least 2, which is strictly larger than $1\vee\alpha^{-1}$ in (1.8). Furthermore, [4, 20] established that $h_{c}(\beta)>0$ for any $\beta>0$, and later [6, 16] obtained the sharp asymptotics for $h_{c}(\beta)$ as $\beta\to 0$. This is called the disorder relevant regime.
@@
$\bullet$ For $\alpha=\frac{1}{2}$, the situation becomes intricate, and it is out of the scope of the Harris criterion. A crucial quantity is $R_{N}:=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}$, the expected number of overlaps for two i.i.d. copies of the renewal process $\tau$. It is shown in [7] that $h_{c}(\beta)>0$ for any $\beta>0$ (and hence the disorder is relevant) if and only if $R_{N}\sim\sum_{n=1}^{N}1/(nL(n)^{2})\to+\infty$. In this scenario, it is called the marginally relevant regime, and [7] provides the sharp asymptotics for $h_{c}(\beta)$ as $\beta\to 0$.
@@
In a nutshell, for $\alpha\geq\frac{1}{2}$ and when the disorder is relevant, we observe significantly different behavior of the pinning model near the critical curve $h=h_{c}(\beta)$ at $(\beta,h_{c}(\beta))=(0,0)$. The disordered system and the pure system belong to two universality classes (cf. [11, Section 1.3]): it is believed that there are different universal scaling limits for the two systems, where neither the disorder distribution nor the details of the renewal process are relevant (as long as some mild conditions hold, e.g., (1.1) holds and the exponent $\alpha$ is fixed).
@@
It is then interesting to study the interpolation of the two universalities. In view of the partition function, as per (1.5), in the localized regime, $Z_{N,\beta}^{\omega,h}$ exhibits exponential growth, determined by the exponent $NF(\beta,h)$. This exponent increases as $N$ tends to infinity and decreases as both $\beta$ and $h$ approach 0. Conversely, in the delocalized regime, $Z_{N,\beta}^{\omega,h}$ demonstrates sub-exponential growth, with $Z_{N,0}^{\omega,0}\equiv 1$. A natural question about interpolating is: if we systematically tune down $\beta:=\beta_{N}\downarrow 0$ and $h:=h_{N}\downarrow 0$ at specific rates as $N\to\infty$, can we ascertain a non-trivial limit for $Z_{N,\beta_{N}}^{\omega,h_{N}}$? The affirmative answer implies the existence of a limit, termed the weak coupling limit, situating the model within the intermediate disorder regime. The inception of exploring the intermediate disorder regime traces back to [1], focusing on the $1+1$ directed polymer model. Subsequent investigations into weak coupling limits include disorder relevant models such as the 2d random field Ising model, the pinning model with $\alpha>1/2$ in [11], and marginally relevant models such as the 2d stochastic heat equation, the $1+2$ directed polymer model, and the pinning model with $\alpha=1/2$ in [12].
@@
For the pinning model with $\alpha>\frac{1}{2}$, a non-trivial weak coupling limit emerges through the choice of $\beta_{N}=\widehat{\beta}{\mathcal{P}}_{1}^{\alpha}(N)$ and $h_{N}=\widehat{h}{\mathcal{P}}_{2}^{\alpha}(N)$ where $\widehat{\beta}>0,\widehat{h}\in{\mathbbm{R}}$, and ${\mathcal{P}}_{1}^{\alpha}(\cdot)$ and ${\mathcal{P}}_{2}^{\alpha}(\cdot)$, both vanishing at infinity, are two regularly varying functions whose exponents depending on $\alpha$ (see [11] for details). However, the scenario becomes more sophisticated in the marginal case $\alpha=\frac{1}{2}$, as illustrated in [12]. The appropriate choice is $h=0$ and $\beta_{N}=\widehat{\beta}/\sqrt{R_{N}}$, where $R_{N}=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}\to+\infty$ as above and is slowly varying at infinity. Furthermore, the constant $\widehat{\beta}$ plays a pivotal role. If $\widehat{\beta}\in(0,1)$, then the weak coupling limit for $Z_{N,\beta_{N}}^{\omega,0}$ is log-normal, characterizing a subcritical phase. For the critical phase $\widehat{\beta}=1$ and the supercritical phase $\widehat{\beta}>1$, $Z_{N,\beta_{N}}^{\omega,0}\to 0$ in probability as $N\to\infty$.
@@
While the behavior of the partition function in the supercritical phase remains mysterious, a recent breakthrough by Caravenna, Sun, and Zygouras in [15] concerning the marginally relevant directed polymer in two dimension sheds light on the critical phase. They viewed the random field of point-to-point partition functions as a process of random measures on ${\mathbbm{R}}^{2}\times{\mathbbm{R}}^{2}$, and established the existence of a universal scaling limit known as the critical 2d stochastic heat flow. Moreover, the limit can be interpreted as the solution of the stochastic heat equation in the critical dimension $d=2$ and in the critical window, which lies beyond existing solution theories via regularity structures by Hairer [25] and paracontrolled distributions by Gubinelli et. al. [24] for subcritical singular SPDEs. Motivated by their work, our study in this paper is centered on the critical phase of the marginally relevant pinning model (Subsection 1.2), as well as a rough volatility model characterised by the critical stochastic Volterra equation (Subsection 1.3).
1.2 The pinning model and weak coupling limit
@@
Instead of the most general setting (1.2), we focus on a specific class of pinning models associated to a random walk $S:=(S_{n})_{n\geq 0}$ on $\mathbb{Z}$ due to technical limitation. To be specific, our proof requires that the renewal process is associated to a Markov process with sharp enough local limit theorems (see Remark 1.4). Noting that the times $S$ visiting 0 give rise to a renewal process, it naturally induces a pinning model. Denote the probability law of $S$ by ${\mathbf{P}}^{x}$ if ${\mathbf{P}}^{x}(S_{0}=x)=1$, indicating that the starting point of $S$ is $x$. The expectation with respect to ${\mathbf{P}}^{x}$ is denoted by ${\mathbf{E}}^{x}$. When $x=0$, we omit the superscripts in $\mathbf{P}$} and $\mathbf{E}$} to lighten the notation. The distribution of the induced renewal process is then given by
$K(n)={\mathbf{P}}(\tau_{1}=n)={\mathbf{P}}(S_{n}=0,S_{k}\neq 0,1\leq k\leq n-1)\quad\text{and}\quad{\mathbf{P}}(n\in\tau)={\mathbf{P}}(S_{n}=0).$ (1.9)
@@
${\mathbf{E}}[S_{1}]=0,\quad{\rm Var}(S_{1})=1,\quad\text{and}\quad{\mathbf{E}}[|S_{1}|^{2+\kappa}]<+\infty~\text{for some}~\kappa>0,$ (1.10)
where $\kappa$ could be arbitrarily small. The assumption on the finite $(2+\kappa)$-th moment is due to technical reasons (see Remark 1.4).
It is known (see [37]) that for a random walk on $\mathbbm{Z}$} with mean 0 and variance $1$, the induced renewal process $\tau$ satisfies $K(n)\sim 1/(\sqrt{2\pi}\hskip 2.0ptn^{3/2})$, which falls within the framework of the marginally relevant pinning model, since $\alpha=\frac{1}{2}$ in (1.4). Furthermore, ${\mathbf{P}}(S_{n}=0)\sim 1/\sqrt{2\pi n}$ by the local limit theorem, and
$R_{N}=\sum_{n=1}^{N}{\mathbf{P}}(n\in\tau)^{2}=\sum_{n=1}^{N}{\mathbf{P}}(S_{n}=0)^{2}\sim\frac{\log N}{2\pi}\to\infty,\quad\text{as~}N\to\infty.$ (1.11)
@@
$\frac{{\mathrm{d}}{\mathbf{P}}_{N,\beta_{N}}^{\omega}}{{\mathrm{d}}{\mathbf{P}}}(S):=\frac{1}{Z_{N,\beta_{N}}^{\omega}}\exp\Big(\sum_{n=1}^{N}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big),$ (1.12)
@@
$Z_{N,\beta_{N}}^{\omega}:={\mathbf{E}}\Big[\exp\Big(\sum_{n=1}^{N}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big)\Big].$ (1.13)
Note that as in [12], here we set the external field $h\equiv 0$ to study the intermediate disorder regime, which is for simplicity and not essential. We are satisfied with this setting and will briefly discuss nonzero choices of $h$ later in Remark 2.4.
Our goal is to study the weak coupling limit of $Z_{N,\beta_{N}}^{\omega}$ at the critical temperature $\beta_{N,c}:=1/\sqrt{R_{N}}\sim\sqrt{2\pi}/\sqrt{\log N}$. Recall that $Z_{N,\beta_{N,c}}^{\omega}$ converges in probability to the trivial limit 0 as $N\to\infty$. The right way to study the weak coupling limit, in the same spirit as [15], is to introduce a random measure related to $Z_{N,\beta_{N,c}}^{\omega}$ and consider its scaling limit. To this end, for given starting and end times $M\leq K$, we introduce a point-to-point pinning partition function
$Z_{M,K}^{\omega,\beta_{N}}:={\mathbf{E}}\Big[\exp\Big(\sum\limits_{n=M}^{K}(\beta_{N}\omega_{n}-\lambda(\beta_{N}))\mathbbm{1}_{\{S_{n}=0\}}\Big)\mathbbm{1}_{\{S_{K}=0\}}\Big|S_{M}=0\Big].$ (1.14)
@@
${\mathcal{Z}}_{N}^{\beta_{N}}({\mathrm{d}}s,{\mathrm{d}}t):=\sqrt{N}Z_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}{\mathrm{d}}s{\mathrm{d}}t,$ (1.15)
@@
$\sqrt{N}{\mathbbm{E}}\big[Z_{\lfloor sN\rfloor,\lfloor tN\rfloor}^{\omega,\beta_{N}}\big]=\sqrt{N}{\mathbf{P}}\big(S_{\lfloor tN\rfloor-\lfloor sN\rfloor}=0\big)\sim\frac{1}{\sqrt{2\pi(t-s)}}.$ (1.16)
We equip the space of locally finite measures on ${\mathbbm{R}}^{2}_{\leq}$ with the topology of vague convergence:
$\mu_{n}\to\mu\quad\Longleftrightarrow\quad\int\phi(s,t)\mu_{n}({\mathrm{d}}s,{\mathrm{d}}t)\overset{\rm d}{\longrightarrow}\int\phi(s,t)\mu({\mathrm{d}}s,{\mathrm{d}}t),\quad\forall\phi\in C_{c}({\mathbbm{R}}^{2}_{\leq}).$ (1.17)
@@
$\zeta_{n}=\zeta_{n}^{(N)}:=e^{\beta_{N}\omega_{n}-\lambda(\beta_{N})}-1.$ (1.18)